Two curves having an equation of x=2√y and y=2√x intersect each


other. Compute the area between the two curves.

1
Expert's answer
2022-06-07T04:53:15-0400

Let's rewrite the function "x=2\sqrt{y}."

"x^2=(2\sqrt{y})^2, x\geq 0, y\geq 0;"

"x^2=4y, x\geq 0, y\geq 0;"

"y=\frac{1}{4}x^2, x\geq 0, y\geq 0."


Let's find the interseptions of the curves:

"\frac{1}{4}x^2=2\sqrt{x},"

"(\frac{1}{4}x^2)^2=(2\sqrt{x})^2,"

"\frac{1}{16}x^4=4x,"

"x^4=64x,"

"x^4-64x=0,"

"x(x^3-64)=0,"

"x=0" or "x^3-64=0,"

"x=0" or "x=4."




"A=\int_0^42\sqrt{x}dx-\int_0^4 \frac{1}{4}x^2dx="


"=2\int_0^4\sqrt{x}dx- \frac{1}{4}\int_0^4x^2dx="


"=2\cdot \frac{2}{3}x\sqrt{x} |_0^4-\frac{1}{4}\frac{x^3}{3}|_0^4="


"=\frac{4}{3}\cdot 4\sqrt{4}-\frac{1}{4}\cdot \frac{4^3}{3}|_0^4=\frac{32}{3}-\frac{16}{3}=\frac{16}{3}=5\frac{1}{3}."


Answer: "5\frac{1}{3}."